OpenStudy (anonymous):

A water holding tank measures 90 m long, 60 m wide, and 9 m deep. Traces of mercury have been found in the tank, with a concentration of 65 mg/L. What is the total mass of mercury in the tank? Answer in units of kg

OpenStudy (nilankshi):

ya........

OpenStudy (anonymous):

the volume of the tank = l*b*h = 60*9*90 =48600 m^3 1 m^3 = 1000 ltrs so 48600 m^3 = 486 * 10^5 ltrs 1litre contains 65 mg= 6.5 gm 486 * 10^5 ltrs contain 6.5 * 486 *10^5 gms =3195 * 10^5 gms = 319500 kgs